Opened 5 years ago

Closed 2 years ago

#533 closed flight processing (fixed)

GB13/08, flight day 217/2014, Montrose Bay

Reported by: knpa Owned by: benj
Priority: alpha 4 medium Milestone: 2014 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by dap)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/GB13_08-2014_217_Montrose_Bay

Data arrived from ARSF via network transfer on 06/08/2014

GB13/08 Scientific objective: Monitoring changes in coastal geomorphology after storm events with LiDAR.

GB13/08 Priority: 8

PI: Mark Cutler

Notes: FW is present. VNIR is x4 spectrally binned.

Sensors:

Camera (18/11/2014)
Fenix (18/11/2014)
LiDAR (18/11/2014)
FW LiDAR (18/11/2014)
OWL (Not requested)

Change History (40)

comment:1 Changed 5 years ago by tec

Started Nav Processing

comment:2 Changed 5 years ago by knpa

  • Description modified (diff)

comment:3 Changed 5 years ago by tec

  • Description modified (diff)
DUDE Basestation Data
Latitude56 27 50.14821
Longitude-2 52 38.33745
El Height57.480m

comment:4 Changed 5 years ago by tec

  • Description modified (diff)

comment:5 Changed 5 years ago by tec

Processing the nav, starting at GPS time 208333. Waiting until tomorrow, for new lever arm values.

comment:6 Changed 5 years ago by tec

Started RCD Processing, converted the Raw files to tifs.

comment:7 Changed 5 years ago by tec

Removed image 05085817100001G4 due to overexposure.

comment:8 Changed 5 years ago by tec

Processing Nav:

Need to get new basestation when smartnet decides to start working again.

Last edited 5 years ago by tec (previous) (diff)

comment:9 Changed 5 years ago by tec

NAV
Using new basestation to try and get a better ambiguity fix

KINT Basestation Data
Latitude57 14 43.34057
Longitude-2 19 43.23940
El Height141.242m
Last edited 5 years ago by tec (previous) (diff)

comment:10 Changed 5 years ago by tec

Nav
Used the old base station in the end. Cut the time to 21160 till 215228. The nav is not perfect over the cross line, it has an quality of 1 for most of the line but there is a bit near the end which is of quality 2.
Nav finishing

Last edited 5 years ago by tec (previous) (diff)

comment:11 Changed 5 years ago by tec

RCD
Delivery created, awaiting DC

comment:12 Changed 5 years ago by dap

Beginning RCD delivery check.

comment:13 Changed 5 years ago by dap

RCD Delivery Check

Complete. The following issues need to be addressed before delivery:

  • The following images have a red tint, which should be commented on in the read me:
    • 84
    • 85
    • 86
    • 87
    • 88
    • 89
    • 102
    • 103
    • 117
    • 118
    • 119
    • 133
    • 136
    • 139
    • 140
    • 141

comment:14 Changed 4 years ago by lah

Started Lidar Processing

comment:15 Changed 4 years ago by lah

Lidar
Processed with variable roll and pitch offsets. Moving onto classification.

Roll Pitch
1 -0.006625 0.000463
2 -0.006625 0.000513
3 -0.006625 0.000513
4 -0.006475 0.000263
5 -0.006475 0.000513
6 -0.006525 0.000513

comment:16 Changed 4 years ago by lah

Lidar
Lidar ready for delivery check. Line 1 FW data was not processed correctly (tried repeating 4 times), so could not be classified correctly. The new script did classify without an error, but the output file could not be read by lasinfo, so the issue seems to be with the header. Reprocessed fine with the windows version of alspp, but there appears to be a slight (0.1m) offset between the 2 las1.2 files using identical settings.

comment:17 Changed 4 years ago by tec

RCD Processing
Ready for delivery.

comment:18 Changed 4 years ago by dac

LiDAR Delivery Check
Starting lidar delivery check.

comment:19 Changed 4 years ago by tec

Hyperspectral Processing
Started processing fenix data.

comment:20 Changed 4 years ago by dac

LiDAR Delivery Check

Delivery check complete. Some minor issues (see below).
Ready to deliver once elevation offsets have been updated in readme.

  • Overlap between 4 -- 6 and 5 --6 is just water, don't think reporting elevation difference makes much sense, would just add note. For 3 -- 6, the difference looks less than 25 cm when only non-water areas are considered.
  • Used full size screenshots in readme (should have used versions created by script with convert -thumbnail 1024) - have corrected.
  • Suggest changing note about data quality to "A large proportion of the flightlines cover water, over which there is a low density of returns. Areas of water have not been classified, although some isolated points have been flagged as class 7 as part of the noise detection algorithm applied." - have done this in readme.

comment:21 Changed 4 years ago by tec

Hyperspectral Processing
APL was moaning about the nav, so renamed the 2014217_ipaspro* files to 2014217_080000*, moved the old files with those names to a different directory. Seems happy now.

comment:22 Changed 4 years ago by tec

Hyperspectral Processing
SCTs

Flightline FENIX
1 0.98
2 0.96
3 0.95
4 -0.05
5 0.92
6 0.90
Last edited 4 years ago by tec (previous) (diff)

comment:23 Changed 4 years ago by tec

Hyperspectral Processing
Finished, ready for delivery check.

comment:24 Changed 4 years ago by lah

'Lidar
Have re-measured overlap between 3 & 6 as 8cm and updated readme, but suggested screenshots don't exist. Not sure what's wrong with script.

comment:25 Changed 4 years ago by lah

Lidar Complete
Recreated readme with thumbnails.

comment:26 Changed 4 years ago by dap

Beginning Hyperspectral Delivery Check

comment:27 Changed 4 years ago by dap

Hyperspectral Delivery Check

Delivery all looks good, just some corrections need to be made in the read me data quality remarks. I've made some corrections to the source tex file but did not run pdflatex on it as there are still some corrections to be made:

  • I changed "final flightline" to "crossline".
  • I changed "is poor" to "are poor".
  • I changed "a high number on" to "high number of"
  • I changed file name of the dem in the example commands section of the read me as check_apl_cmd returned some errors stating that the dem file didn't exist (it was expecting a dem with a different name). I've also changed the command in the read me config file.

I didn't make corrections to remarks on the underflows as I wasn't sure whether these needed to be corrected. Upon inspection of the level1b bil files, I found the following noisy bands:

Flightline Noisy bands
01242-256
320-371
390-415
420-440
02242-256
320-359
390-415
420-447
03321-370
420-447
04320-362
420-447
05320-350
416-447
06245-265
315-447

comment:28 Changed 4 years ago by tec

Hyperspectral Processing
Implemented changes, slightly generalised bands so that I could group lines 1,2 and 3,4,5

Ready to be checked again.

Last edited 4 years ago by tec (previous) (diff)

comment:29 Changed 4 years ago by dap

Hyperspectral Delivery Check

Checked corrections to read me document, hyperspectral data ready for delivery.

comment:30 Changed 4 years ago by dap

PI has been sent pre vector delivery email, awaiting his reply before delivering the data.

comment:31 Changed 4 years ago by dap

Zipping mapped files for copying delivery to hard drive.

comment:32 Changed 4 years ago by dap

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

Renamed DEM and hdr so that file names comply with naming convention.

LiDAR, Hyperspectral and Camera data delivered to Mark Cutler at the University of Dundee, 18/11/2014

comment:33 Changed 4 years ago by mark1

Fenix
Fenix data needs reprocessing due to data being binned spectrally where processing used an incorrect algorithm. Also, SWIR integration time was recorded incorrectly when data collected. Have changed tint2 values from 26.2 to 2.15 as per Specim's suggestion.

Data are being reprocessed.

comment:34 Changed 4 years ago by dap

Fenix Delivery Check Take 2

Spectral profile of vegetation looks as expected.

Just a couple of other minor points to comment on:

  • Cross line looks as though it still has some wobbles in the line (particularly over the water) but the data quality remarks state that the navigation data is of a poor quality, so not much we can do about this I think.
  • The 4th link on page 8 in the read me (https://lpdaac.usgs.gov/products/aster_policies) returns a 404 HTTP Error Code.
  • xml files have real paths pointing to molucca2 - will this be a problem if we run the example commands from there when(/if) the deliveries are moved to visuyan2?

comment:35 Changed 4 years ago by dap

Fenix

Fixed the broken link in the read me and added acknowledgement for the ASTER DEM. Zipping mapped files.

comment:36 Changed 4 years ago by dap

Fenix

Mapped files have been zipped, Fenix delivery on molucca2 now ready for delivery.

comment:37 Changed 3 years ago by gej

Reprocessed Fenix Delivered
Delivered to Mark Cutler School of the Environment, University of Dundee 27th August 2015

comment:38 Changed 3 years ago by stgo

Archiving

Starting archiving.

comment:39 Changed 3 years ago by dac

Archiving

Processed data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2014/GB13_08/GB13_08-2014_217_Montrose_Bay). Raw data not available. Started uploading again.

comment:40 Changed 2 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Raw data now available from NEODC. Marking as closed.

Note: See TracTickets for help on using tickets.